OpenStudy (anonymous):

A store had 175 cell phones in the month of January. Every month, 10% of the cell phones were sold and 10 new cell phones were stocked in the store. Which recursive function best represents the number of cell phones in the store f(n) after n months?

OpenStudy (anonymous):

(A) f(n) = 175 - 0.9 × f(n - 1) + 10, f(0) = 175, n > 0 (B) f(n) = 0.1 × f(n - 1) + 10, f(0) = 175, n > 0 (C) f(n) = 175 + 0.9 × f(n - 1) + 10, f(0) = 175, n > 0 (D) f(n) = 0.9 × f(n - 1) + 10, f(0) = 175, n > 0

OpenStudy (anonymous):

(A) f(n) = 175 - 0.9 * f(n - 1) + 10, f(0) = 175, n > 0 (B) f(n) = 0.1 * f(n - 1) + 10, f(0) = 175, n > 0 (C) f(n) = 175 + 0.9 * f(n - 1) + 10, f(0) = 175, n > 0 (D) f(n) = 0.9 * f(n - 1) + 10, f(0) = 175, n > 0